OpenStudy (anonymous):

consider the arithmetic sequence -15, -11, -7, -3 write an equation for the 9nth term find the 25th term

OpenStudy (anonymous):

The common difference is d=-4 The nth term of an arithmetic sequence is given by\[a (n)=a_{1}+(n-1)d\]In this case you have\[a(9)=-15+(9-1)(-4)=-15-32=-47\]

OpenStudy (anonymous):

The common difference if what you would add to each term to get the next term.

OpenStudy (anonymous):

whbout the 25th term?

OpenStudy (anonymous):

\[a(25)=-15+(25-1)(-4)=-15-96=-111\]